OpenStudy (anonymous):

find the equation of a straight line that has a slope of m=2/3 and is passing thru the point of -2,-4

OpenStudy (anonymous):

(y-y1) = m(x-x1) where your (-2,-4) = (x1,y1) and m=2\3 so that becomes, (y-(-4)) = 2\3(x-(-2)) (y+4) = 2\3(x+2) 3(y+4) = 2(x+2) distribute 3 and 2, 3y + 12 = 2x + 4 3y - 2x +12 - 4 = 0 ans.. 3y - 2x +8 = 0 ^_^